I'm supposed to solve and check this fractional equation.


(x-5 over x) + (3 over x) = (2 over 3)

Answer by stanbon(48551)

(x-5 over x) + (3 over x) = (2 over 3)
---------
(x-5)/x + 3/x = 2/3
Add the fraction on the left side of the equal sign.
(x-2)/x=2/3
Cross-multiply.
3(x-2)=2x
3x-2x=6
x=6
=========
Cheers,
Stan H.